### v1

# AI-Augmented Thyroid Scintigraphy for Robust Classification

Abstract
Thyroid scintigraphy is a key imaging modality for diagnosing thyroid disorders. Deep learning models for thyroid scintigraphy classification often face challenges due to limited and imbalanced datasets, leading to suboptimal generalization. In this study, we investigate the effectiveness of different data augmentation techniques—Stable Diffusion (SD), Flow Matching (FM), and Conventional Augmentation (CA)—to enhance the performance of a ResNet18 classifier for thyroid condition classification. Our results showed that FM-based augmentation consistently outperforms SD-based approaches, particularly when combined with original (O) data and CA (O+FM+CA), achieving both high accuracy and fair classification across Diffuse Goiter (DG), Nodular Goiter (NG), Normal (NL), and Thyroiditis (TI) cases. The Wilcoxon statistical analysis further validated the superiority of O+FM and its variants (O+FM+CA) over SD-based augmentations in most scenarios. These findings highlight the potential of FM-based augmentation as a superior approach for generating high-quality synthetic thyroid scintigraphy images and improving model generalization in medical image classification.
Keywords:
Thyroid Scintigraphy Image synthesis Augmentation Diffusion Stable diffusion Flow matching
1 Introduction
Thyroid diseases are among the most common endocrine disorders, affecting millions of subjects worldwide [17]. Early and accurate diagnosis is essential for effective treatment and better patient outcomes. Physicians rely on various imaging techniques, including ultrasound (US), computed tomography (CT), magnetic resonance imaging (MRI), and thyroid scintigraphy (gamma scan), along with laboratory tests, to assess thyroid conditions. Although US is widely used to evaluate nodules, its precision depends on operator’s expertise. CT and MRI are able to assess structural abnormalities but often yield nonspecific results, especially in cases like Graves’ disease [6]. By contrast, thyroid scintigraphy (using the 99mTc-pertechnetate radiopharmaceutical) provides crucial insight into the structure and function of the thyroid gland. However, interpreting these images can be subjective, time-consuming, and prone to variability among experts. These challenges highlight the need for improved, automated approaches to enhance diagnostic accuracy and efficiency [7]. Artificial intelligence (AI) has shown significant potential in the medical domain, particularly in disease diagnosis, treatment guidance, and personalized patient care [2]. However, a major challenge in developing deep learning (DL) models for medical applications is data scarcity. Large datasets are crucial for training robust models, yet collecting sufficient data can be difficult due to privacy concerns, high costs, and logistical constraints [19]. Conventional augmentation (CA) techniques, such as rotation, flipping, shifting, scaling, etc. help improve model generalization by creating variations of existing data [12]. However, these methods alone are often insufficient to fully address data limitations, highlighting the need for more advanced strategies [22]. Recent studies have investigated advanced augmentation techniques to overcome the challenge of limited medical imaging datasets. While Generative Adversarial Networks (GANs) [8] and Variational Autoencoders (VAEs) [13] have shown success, diffusion-based models [11] demonstrate superior performance in image synthesis, producing highly realistic augmented images [9, 22, 1]. This is the first study to implement a comprehensive augmentation approach on thyroid scintigraphy images using diffusion-based algorithms, including Denoising Diffusion Probabilistic Models (DDPM) [11] and Flow-Matching (FM) [15] to address the challenge of limited medical imaging data. To assess the effectiveness of the generated images, we incorporate them into the training process of a classification model and evaluate their performance on an external dataset. Contribution: Our key contributions are as follows: 1) A novel application of diffusion-based models for augmenting thyroid scintigraphy images to address data scarcity. 2) Prompt generation from physicians’ reports to maximize the use of available information in the image synthesis process. 3) A demonstration of the effectiveness of diffusion-based augmentation in enhancing classification performance for thyroid scintigraphy imaging.
2 Related-works
Zhang et al. [22] investigated SinDDM [14], a single-image denoising diffusion model, to augment lung ultrasound data. They also introduced FewDDM, an extension trained on limited samples, which outperformed single-image GANs in generating high-quality synthetic images. Augmenting with SinDDM notably improved pathology classification, especially for minority classes. Despite generating less detailed images, FewDDM surpassed SinDDM and SinGAN in downstream performance by capturing local structural variations. The study highlighted that combining synthetic and CA techniques yielded the best classification results. Hajianfar et al. [9] investigated the effectiveness of stable diffusion (SD) [18] as an advanced augmentation method in enhancing deep learning models for classifying scintigraphic thyroid images. They used reports from physicians without specific cleaning as prompts in the augmentation process. The generated images, combined with original (O) data and CA, were used to train a classifier. The results demonstrated that models trained with synthetic data achieved consistently better performance. Balla et al. [3] explored strategies to address data scarcity in musculoskeletal US for osteoarthritis detection. They used CA with diffusion-based image synthesis and the results showed that synthetic images generated through diffusion models retained anatomical fidelity and improved model generalization diagnostic accuracy, while CA sometimes hindered performance highlighting the potential of using synthetic images. Akrout et al. [1] advances data augmentation by leveraging text-to-image diffusion models to enhance a macroscopic skin disease dataset. By using text prompts, they gain fine-grained control over the image generation process. The results show that this generative augmentation approach maintains classification accuracy even when trained on a fully synthetic dataset. FM [15], as a novel, more robust, and memory-efficient method for image synthesis has shown superiority over GANs and DDPMs. While it has not been widely used in the medical domain, it has demonstrated clear advantages. In this study, we employed various strategies for advanced augmentation. Specifically, we use image masks as conditions for both DDPM and FM, and we incorporate physicians’ reports as prompts for DDPM to maximize the available information for augmentation. By leveraging FM, we aim to improve both the efficiency and quality of image synthesis, making it a key component of our approach.
3 Methodology
We aimed to find the best augmentation method to enhance the classification performance. Conventional methods [20] can generate samples that belong to completely different classes [12]. Therefore, we need a method to learn the distribution of the dataset images and then draw samples from it. GANs[8], Variational Autoencoders (VAEs) [13], DDPMs [11] and FMs [15] are some examples. Among these algorithms, DDPMs and FMs have exhibited superior performance [4]. Hence, we only consider these two approaches.
Stable Diffusion: SD is a type of Latent Diffusion Model (LDM)[18], which belongs to the DDPM family but operates in a lower-dimensional latent space to improve efficiency. LDMs first encode the image into a compact latent representation using a pre-trained VAE. The diffusion process then operates in this latent space, making it computationally efficient. For the diffusion process, there are two phases: the Forward and Reverse Processes. Let the target image be denoted as $\mathbf{x}_{0}$. In the forward process, we gradually add Gaussian noise to the sample in a Markovian manner: $\mathbf{x}_{t}=\sqrt{1-\beta_{t}}\mathbf{x}_{t-1}+\sqrt{\beta_{t}}\mathbf{v}_{t},\quad\mathbf{v}_{t}\sim\mathcal{N}(0,I)$. The coefficients $\sqrt{1-\beta_{t}}$ and $\sqrt{\beta_{t}}$ control the transition, ensuring a gradual corruption of the data while maintaining variance stability. After a sufficient number of steps ($T$), the sample $\mathbf{x}_{T}$ follows a standard Gaussian distribution, i.e., $\mathbf{x}_{T}\sim\mathcal{N}(0,I)$. To generate new samples, we approximate the reverse diffusion process. Starting from $\mathbf{x}_{T}\sim\mathcal{N}(0,I)$, we iteratively sample from the conditional distribution, i.e., $\mathbf{x}_{t-1}\sim p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t}).$ Since $q(\mathbf{x}_{t-1}|\mathbf{x}_{t})$ is intractable, we assume it follows a Gaussian distribution and train a neural network (e.g., a U-Net) to estimate the necessary parameters for sampling. By iterating this denoising process from $T$ down to 00, we obtain a generated sample $\hat{\mathbf{x}}_{0}$. Since we are using the SD model, we can incorporate additional conditioning information into the reverse process. This is achieved by modifying the reverse process to be conditional on auxiliary inputs such as text prompts or images. The new conditional distribution is given by: $\mathbf{x}_{t-1}\sim p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},C).$ where $C$ represents the selected condition, which can be $P$ (a prompt), $M$ (a mask), or $y$ (a given image). To enable conditioning, SD trains the noise prediction model $p_{\theta}$ to take both $x_{t}$ and $C$ as inputs, ensuring that the generated sample aligns with the provided condition.
Flow matching: FM provides an alternative to diffusion models by directly learning a continuous-time velocity field that defines a near-optimal transport between the source and target distributions. FM defines a straight-line (or nearly straight) transformation between samples from the data distribution and a known prior. This makes sampling more efficient compared to traditional diffusion-based approaches. Let $\mathbf{x}_{0}\sim p_{0}(x)$ and $\mathbf{x}_{1}\sim p_{1}(x)$ represent two distributions, where $\mathbf{x}_{0}$ is the source distribution (e.g., real data) and $\mathbf{x}_{1}$ is the target distribution (e.g., noise). FM constructs a continuous interpolation between these two distributions as: $\mathbf{x}_{t}=t\mathbf{x}_{1}+(1-t)\mathbf{x}_{0},\quad t\in[0,1].$ This formulation defines a linear transport path from $\mathbf{x}_{0}$ to $\mathbf{x}_{1}$. The goal is to learn a velocity field $\mathbf{v}_{\theta}(\mathbf{x}_{t},t)$ that describes the optimal transport direction at each time step. Ideally, this velocity field should satisfy: $v_{\theta}(\mathbf{x}_{t},t)=\mathbf{x}_{1}-\mathbf{x}_{0}.$ To ensure that the learned velocity field $\mathbf{v}_{\theta}(\mathbf{x}_{t},t)$ correctly follows the transport direction, we minimize the FM loss: $\mathcal{L}=\mathbb{E}_{\mathbf{x}_{0},\mathbf{x}_{1}}\left[|(\mathbf{x}_{1}-\mathbf{x}_{0})-v_{\theta}(\mathbf{x}_{t},t)|_{2}^{2}\right].$ During inference, novel samples can be generated by solving the learned ordinary differential equation (ODE) defined by the velocity field: $\frac{d\mathbf{x}_{t}}{dt}=v_{\theta}(\mathbf{x}_{t},t).$ This ODE governs the smooth transport from $\mathbf{x}_{0}$ to $\mathbf{x}_{1}$. Unlike diffusion models, which require many discretized steps for effective denoising, FM provides a single-step or low-step approximation to recover the target distribution.
4 Experiments
Data collection: Table 1 provides an overview of the studied cases collected from nine centers using eight different imaging systems. The dataset covers a broad age range with a mean age of 44.71 ± 17.66 years. The gender distribution includes 771 males (26%) and 2,183 females (74%). The cases are classified into four categories: Diffuse Goiter (DG), Nodular Goiter (NG), Thyroiditis (TI), and Normal (NL), totaling 2,954 cases.
Table 1: Data information and distribution across different centers.
Center Age M/F DG/NG/TI/NL Total Manufacturer Model
Training Dataset
A NA 129/303 100/203/81/48 432 ADAC GENESYS
B 46.12 ± 15.26 77/166 63/110/47/23 243 SIEMENS IP2 (ECAM1028)
C 45.08 ± 14.84 137/511 215/317/88/28 648 SIEMENS IP2 (ECAM1028)
D 43.81 ± 22.26 150/448 145/199/134/120 598 Mediso AnyScan
E 47.04 ± 16.10 70/249 89/121/60/49 319 SIEMENS IP1 (ECAM10482)
F 41.42 ± 14.75 91/224 176/61/49/29 315 GE Discovery NM 630
External Dataset
G 42.90 ± 12.00 21/50 20/21/24/6 71 MiE SCINTRON
H 46.78 ± 15.64 46/96 61/42/31/8 142 SIEMENS Encore 2 (SYMBIA1071)
I 41.30 ± 15.20 50/136 46/35/60/45 186 GE INFINIA
Total 44.71 ± 17.66 771/2183 915/1109/574/356 2954
M: Male, F: Female, DG: Diffuse Goiter, NG: Nodular Goiter, TI: Thyroiditis, NL: Normal
Preprocessing: An experienced nuclear medicine physician manually segmented the thyroid region from the scintigraphy images using the manual contouring tool in ITK-Snap software [21]. Among the 2,954 images, 319 (all from center E) had a resolution of 256 × 256, while the rest were 128 × 128. Images and masks from center 5 were resampled to 128 × 128 using BSpline and nearest neighbor interpolation, respectively. Data usage for this study was approved by the research ethics board of each participating institution. In this study, we use physicians’ case reports in synthesizing images, so consistency is crucial due to varying styles and approaches across different centers. First, we analyzed all reports using GPT-4 Turbo [16] and generated questions to extract the most relevant information. These questions were then reviewed and refined by an experienced nuclear medicine physician. Finally, we used the revised questions to gather consistent information using GPT-4 Turbo and create prompts under 77 tokens to feed SD (Table 2). Additionally, for each center, an experienced technician randomly reviewed 50 cases of the generated prompts.
Table 2: Structured information extraction from thyroid scan reports
[…]
Your task is to analyze a thyroid scan report and extract structured information."
Given the following thyroid scan report: "{report}"
Answer the following questions in a structured JSON format:
"Class": "DG, NG, TI, NL",
"thyroid_function_classification": "Hypofunction / Hyperfunction / Normal / Indeterminate",
"radiotracer_uptake_pattern": "Homogeneous / Inhomogeneous / Focal / Diffuse",
"has_nodules": "Yes / No",
"multinodular_goiter": "Yes / No",
"nodule_type": "Hot / Cold / Not specified",
"thyroid_size": "Normal / Mildly enlarged / Significantly enlarged / Atrophic",
"diffuse_enlargement": "Yes / No"
4.1 Training set-up
We trained our models using data from six centers (A–F) and evaluated the classifier on an external dataset from three centers (G–I) for classification evaluation. Five augmentation methods were employed, including CA, three variations of SD, and FM. For each augmentation method, 1,000 images were generated per class.
Conventional augmentation: We applied a randomized transformation pipeline that included rotation (±15°), horizontal flipping, translation (±10%), scaling (0.8–1.2×), and Gaussian noise addition ($\sigma$ = 0.001–0.01).
Stable diffusion augmentation: We used a fine-tuning setup with mixed precision (fp16), a resolution of 128×128, exponential moving average, gradient accumulation with four steps, and a batch size of 1, optimizing for 50,000 steps at a learning rate of 1e-5. During inference, we evaluated three approaches: 1) image and prompt to image (SD1), 2) prompt to image (SD2), and 3) mask and prompt to image (SD3).
Flow matching augmentation: For FM, the model was optimized using the Adam optimizer with a learning rate of 1e-4 for 200 epochs. Our approach leverages FM with optimal transport to align predicted data flows with the target distribution. Class conditioning is achieved by incorporating a one-hot encoded vector via cross-attention, while mask conditioning is implemented through a parallel control network integrated via residual connections. During inference, we assessed guided generation using a combination of mask and class conditioning.
This resulted in 15 distinct training strategies for the ResNet18 classifier: 1) O, 2) CA, 3) SD1, 4) SD2, 5) SD3, 6) FM, 7) O+CA, 8) O+SD1, 9) O+SD2, 10) O+SD3, 11) O+FM, 12) O+CA+SD1, 13) O+CA+SD2, 14) O+CA+SD3, and 15) O+CA+FM.
ResNet18: We trained a ResNet18 classifier, initializing it with ImageNet1K pre-trained weights. The first convolutional layer was modified to retain a 3×3 kernel, and the fully connected layer was replaced with a dropout layer (0.2) followed by a linear layer matching the number of classes. The model was trained for 300 epochs using the Adam optimizer (learning rate = 1e-4, weight decay = 1e-5) with a cross-entropy loss function. A learning rate scheduler (ReduceLROnPlateau) was applied, reducing the rate by a factor of 0.8 if validation loss plateaued for 10 epochs, with a minimum learning rate of 2e-5. Furthermore, the training and validation split was set at a 9:1 ratio with stratification.
4.2 Evaluation metrics
Augmentation metrics: Fréchet Inception Distance (FID) [10] and Kernel Inception Distance (KID) [5] metrics have been used for class-wise and overall comparisons between generated and original images. In both cases, 200 images per class were selected from each dataset for evaluation.
Classification metrics: The classification performance was evaluated on an external dataset using metrics, including precision, recall, F1-score, accuracy, and the area under the Receiver Operating Characteristic curve (ROC AUC). Given that it was a multiclass task, we applied various averaging techniques encompassing micro, macro, and weighted to provide a thorough evaluation across all classes.
Statistical method: We compared different strategies using bootstrapping with 1,000 repetitions and sampling with replacement. Accuracy distributions were analyzed, and pairwise comparisons were made using the Wilcoxon rank sum test, considering p-values < 0.5 as statistically significant.
GradCam: We used Gradient-weighted Class Activation Mapping (GradCAM) on the trained ResNet18 model to generate heatmaps, visualizing the image regions that influenced the classifier’s decisions.
5 Results and analysis
Figure 1 presents one sample per class from the original dataset alongside examples generated by different augmentation methods used in this study. Additionally, it includes samples from the external dataset with their corresponding Grad-CAM visualizations using the O+FM+CA model, highlighting the model’s focus during prediction. The analysis of FID and KID metrics (Table 3) reveals significant performance differences. SD3 performs the worst with the highest FID (17.99) and KID Mean (33.59), indicating poor image quality and large distribution shifts. SD1 and SD2 perform better, with SD2 achieving the lowest FID (3.88), though its TI class shows a high KID Mean (30.02). FM outperforms all, with the best FID (0.66) and minimal KID Mean, indicating its generated images closely match real data. Table 4 and Table 5 highlight the impact of different model configurations on classification performance. The inclusion of FM and CA components consistently improves performance across all metrics, with O+FM+CA and O+FM achieving the highest F1-scores. Notably, O+FM provides the best balance between precision and recall, while FM alone offers strong precision but slightly lower recall. In terms of accuracy and AUC, O+FM outperforms other models, reaching an accuracy of 0.78 and the highest AUC values. SD3, which exhibits low recall and AUC, lags behind other configurations. Additionally, the Wilcoxon rank sum test (Figure 2) confirms that O+FM and O+FM+CA consistently outperform other models, while SD1, which incorporates both prompt and image during inference, outperforms SD2 and SD3. SD2, which lacks image guidance, and SD3, which struggles with masked inputs, show suboptimal results.
Table 3: Comparison of FID and KID Mean (Std < 1e-5 for all) for SD and FM Methods.
Class SD1 SD2 SD3 FM
FID $\downarrow$ KID $\downarrow$ FID $\downarrow$ KID $\downarrow$ FID $\downarrow$ KID $\downarrow$ FID $\downarrow$ KID $\downarrow$
DG 4.75 6.77 2.94 2.31 10.30 15.40 0.96 0.80
NG 6.22 10.28 3.27 2.66 22.07 38.34 0.74 0.95
NL 4.56 7.14 1.75 2.05 18.68 36.74 0.85 2.08
TI 2.73 3.01 9.24 30.02 22.37 50.50 1.97 3.39
Overall 4.17 4.99 3.88 2.61 17.99 33.59 0.66 0.83
Figure 1: Examples of original and augmented images for each class using different methods. Grad-CAM visualizations from the O+FM+CA model are also shown for external dataset samples, highlighting the model’s focus during prediction.
Table 4: Class-wise performance metrics of the classification. (Support: DG: 130, NG: 95, NL:74, and TI: 100)
Method Precision Recall F1-score
(DG/NG/NL/TI) (DG/NG/NL/TI) (DG/NG/NL/TI)
O 0.70/0.65/0.53/0.96 0.72/0.86/0.31/0.91 0.71/0.74/0.39/0.93
CA 0.75/0.70/0.55/0.92 0.62/0.77/0.62/0.97 0.68/0.73/0.58/0.95
SD1 0.80/0.74/0.53/0.92 0.72/0.71/0.59/0.98 0.76/0.72/0.56/0.95
SD2 0.76/0.57/0.45/0.95 0.70/0.68/0.53/0.74 0.73/0.62/0.49/0.83
SD3 0.88/0.66/0.42/0.58 0.44/0.35/0.64/1.00 0.58/0.46/0.50/0.74
FM 0.70/0.54/0.48/0.96 0.67/0.64/0.54/0.76 0.69/0.59/0.51/0.85
O+CA 0.73/0.59/0.61/0.99 0.72/0.83/0.41/0.89 0.72/0.69/0.49/0.94
O+SD1 0.74/0.67/0.55/0.96 0.70/0.85/0.46/0.89 0.72/0.75/0.50/0.92
O+SD2 0.78/0.60/0.61/0.98 0.67/0.87/0.49/0.89 0.72/0.71/0.54/0.93
O+SD3 0.76/0.62/0.59/0.95 0.66/0.87/0.47/0.89 0.71/0.73/0.53/0.92
O+FM 0.74/0.74/0.68/0.93 0.81/0.76/0.54/0.95 0.77/0.75/0.60/0.94
O+SD1+CA 0.75/0.64/0.54/0.96 0.66/0.82/0.50/0.91 0.70/0.72/0.52/0.93
O+SD2+CA 0.74/0.65/0.62/0.94 0.66/0.82/0.54/0.91 0.70/0.73/0.58/0.92
O+SD3+CA 0.72/0.63/0.58/0.95 0.63/0.82/0.53/0.90 0.67/0.72/0.55/0.92
O+FM+CA 0.79/0.68/0.63/0.97 0.73/0.84/0.58/0.90 0.76/0.75/0.61/0.93
Table 5: Averaging performance metrics of the classification. (Support: DG+NG+NL+TI=399)
Method AUC Precision Recall F1-score
(mic/mac/wei) (mic/mac/wei) (mic/mac/wei) (mic/mac/wei)
O 0.92/0.91/0.91 0.73/0.71/0.72 0.73/0.70/0.73 0.73/0.69/0.71
CA 0.93/0.92/0.92 0.74/0.73/0.75 0.74/0.74/0.74 0.74/0.74/0.74
SD1 0.93/0.92/0.92 0.76/0.74/0.76 0.76/0.75/0.76 0.76/0.75/0.76
SD2 0.88/0.87/0.88 0.67/0.68/0.70 0.67/0.66/0.67 0.67/0.67/0.68
SD3 0.82/0.86/0.86 0.59/0.63/0.67 0.59/0.61/0.59 0.59/0.57/0.58
FM 0.89/0.88/0.88 0.66/0.67/0.69 0.66/0.65/0.66 0.66/0.66/0.67
O+CA 0.92/0.92/0.92 0.73/0.73/0.74 0.73/0.71/0.73 0.73/0.71/0.73
O+SD1 0.92/0.92/0.92 0.74/0.73/0.74 0.74/0.73/0.74 0.74/0.72/0.74
O+SD2 0.92/0.92/0.92 0.74/0.74/0.76 0.74/0.73/0.74 0.74/0.73/0.74
O+SD3 0.92/0.92/0.92 0.73/0.73/0.74 0.73/0.72/0.73 0.73/0.72/0.73
O+FM 0.95/0.93/0.94 0.78/0.77/0.78 0.78/0.76/0.78 0.78/0.77/0.78
O+SD1+CA 0.93/0.92/0.93 0.73/0.72/0.74 0.73/0.72/0.73 0.73/0.72/0.73
O+SD2+CA 0.92/0.92/0.92 0.74/0.73/0.74 0.74/0.73/0.74 0.74/0.73/0.74
O+SD3+CA 0.92/0.91/0.92 0.72/0.72/0.73 0.72/0.72/0.72 0.72/0.72/0.72
O+FM+CA 0.93/0.92/0.92 0.77/0.77/0.78 0.77/0.76/0.77 0.77/0.76/0.77
Figure 2: Pairwise model comparison via the Wilcoxon signed-rank test. Each cell compares two models: Blue (row model significantly better), Peach (worse), and Red (no significant difference).
6 Conclusion
This study examined the impact of various augmentation strategies, including SD-, FM-, and CA-based methods, on thyroid classification using ResNet18. Our results show that FM-based augmentation, particularly when combined with the original dataset (O+FM) or CA (O+FM+CA), consistently outperformed SD-based approaches. FM enables smoother, more controlled image transformations, preserving key structural and intensity details crucial for classification.
{credits}
6.0.1 Acknowledgements
This work was supported by the Natural Sciences and Engineering Research Council of Canada (NSERC) Discovery Grant RGPIN-2019-06467.
[…]
The authors have no competing interests to declare.